40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 31 20:06:22 2013, Evan, Update, Locking, PRM/PR2 cavity 
    Reply  Thu Jan 31 23:48:45 2013, Koji, Update, Locking, PRM/PR2 cavity 
       Reply  Fri Feb 1 09:33:11 2013, Jamie, Update, Locking, PRM/PR2 cavity 
          Reply  Fri Feb 1 15:12:53 2013, Jenne, Update, Locking, PRM/PR2 cavity 
             Reply  Fri Feb 1 19:55:33 2013, Jenne, Update, Locking, PRM/PR2 cavity 
                Reply  Fri Feb 1 23:12:42 2013, Koji, Update, Locking, PRM/PR2 cavity 
                Reply  Fri Feb 1 23:52:59 2013, rana, Update, Locking, PRM/PR2 cavity 
Message ID: 7987     Entry time: Fri Feb 1 23:12:42 2013     In reply to: 7986
Author: Koji 
Type: Update 
Category: Locking 
Subject: PRM/PR2 cavity 

During the scanning we were riddled by the fact the PDH error and the transmission peaks do not happen simultaneously.

After a little investigation, it was found that "LP100^2" filter is left on in the POPDC filter.

Moreover, it was also found that the whitening filter switches for the POPDC does not switch the analog counterpart.

These were the culprit why we never saw accidental hitting of the max transmission by the peaks when the cavity was not locked.

 

I know that the most of the whitening filter in the RF paths were checked before (by Keiko?), but the similar failure still exists in the POX path.
We should check for the whitening filters in the DC path as well and fix everything at once. I can offer assistance on the fixing part.

ELOG V3.1.3-